Determine the molarity of a solution in mol/L that is prepared by completely dissolving 1.23 g of solid SrF2 (Mm=125.616g/m) in 417g of water.


1
Expert's answer
2021-09-18T23:58:14-0400

Since we know that M=n/V, we use the initial data we have and then we proceed to calculate M:


"M=\cfrac{1.23\,g\,SrF_2}{417\,g\,H_2O}\times \cfrac{1\,mol\,SrF_2}{125.616\,g\,SrF_2}\times \cfrac{1\,g\,H_2O}{0.001\,L\,H_2O} 
\\ \text{ }
\\ \implies M=0.0235 \,mol/L"


In conclusion, the molarity of the SrF2 solution is M = 0.0235 mol/L.


Reference

  • Chang, R., & Goldsby, K. A. (2010). Chemistry. Chemistry, 10th ed.; McGraw-Hill Education: New York, NY, USA.